On Monday, Bill drove to work in 50 minutes. On Tuesday, he averaged 12 miles per hour more, and it took him 9 minutes less to get to work. how far does he travel to work?
Answer by stanbon(48551)

Bill drove to work in 50 minutes. On Tuesday, he averaged 12 miles per hour more, and it took him 9 minutes less to get to work. how far does he travel to work?
1st Day DATA:
time= 50 min ; rate = x mi/min ; distance=rt = 50x
2nd Day DATA:
time=41 min; rate=(x+(12/60)) mi/min; distance= rt = 41(x+(1/5)=41x+8.2 mi/min
EQUATION:
distance = distance
50x = 41x+8.2
9x = 8.2
x= 0.91111 mi/min
50x=45.56 miles
Cheers,
Stan H.